What is the volume of NO2 gas required to dissolve in 500 cm3 of water to give a 0.100 M
HNO3 solution at room temperature and pressure ?
1
Expert's answer
2020-05-19T08:58:28-0400

2NO2+H2O=HNO3+HNO2 ;

PV=nRT ideal gas law;

V(NO2)=?

1) n(HNO3)=0.100*500/1000=0.05 mol

2) 2 mol(NO2) ____________ 1 mol (HNO3)

X mol _______________ 0.05 mol

3) X=0.05*2/1=0.1 mol NO2

4) V(NO2)=nRT/P=0.1*8.31*273/101.325=2.24 L